find the length of the short leg of a right triangle if the longer leg is 11 feet more than the shorter leg and the hypotenuse is 11 feet less than twice the shorter leg
Answer by nerdybill(7384)

find the length of the short leg of a right triangle if the longer leg is 11 feet more than the shorter leg and the hypotenuse is 11 feet less than twice the shorter leg
.
Let x = length of shorter leg
then
x+11 = longer leg
2x-11 = hypotenuse
.
Applying Pythagorean's theorem:
x^2 + (x+11)^2 = (2x-11)^2
x^2 + (x+11)(x+11) = (2x-11)(2x-11)
x^2 + (x^2+22x+121) = 4x^2-44x+121
2x^2+22x+121 = 4x^2-44x+121
22x+121 = 2x^2-44x+121
121 = 2x^2-66x+121
0 = 2x^2-66x
0 = 2x(x-33)
x = {0, 33}
throw out the zero (extraneous) leaving
x = 33 feet